OpenStudy (anonymous):

4x^3/3 times (9/x)^3

OpenStudy (anonymous):

4x^3/3 * 9/x^3 4/3 * 9 4*3 = 12 <=======

OpenStudy (anonymous):

the ^3 goes to the 9 and the x

OpenStudy (anonymous):

\[\frac{4x^3}{3}*(\frac{9}{x})^3\]?

OpenStudy (anonymous):

yes!

OpenStudy (bahrom7893):

That is: 4*729/3=972

OpenStudy (anonymous):

ah, in that case it is a bit bigger :P 972

OpenStudy (anonymous):

\[\frac{(4x^3)*9^3}{3*x^3}\]

OpenStudy (anonymous):

\[\frac{2.916x^2}{3x^3}=972

OpenStudy (anonymous):

thanks everyone!

OpenStudy (anonymous):

\[\frac{2.916x^3}{3x^3}= 972\]